OpenStudy (anonymous):

What interest rate, compounded annually, is equivalent to a 3% rate compounded continuously? Round to two decimal places.

OpenStudy (anonymous):

I know that annually is the equation P=Poa^t and continuously the equation is P=Poe^kt. Can I make Poa^t=Poe^kt? Even then, I feel like I have nothing to go by in terms of solving it.

OpenStudy (anonymous):

A/P=(1+r) = e^0.03 using t = 1 for both e^0.03 = 1.0304

OpenStudy (anonymous):

That's what I put on my test, and I got it wrong. I have a similar problem I found: "What interest rate, compounded annually, is equivalent to a 7% rate compounded continuously? Round to two decimal places." The answer to that one in the answer key was 7.25%